OpenStudy (skullcandy229):

The area of a rectangle is 4y + 10. A. Represent the area of the rectangle in factored form. B. Use the factored form to give one possible set of dimensions for this rectangle.

OpenStudy (alexandervonhumboldt2):

4y+10=2*2y+2*5=2*(2y+5)

OpenStudy (tkhunny):

That's one way to do it. There are infinitely many other ways to do it. Give us another!